Circular No. 7946

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 57P/DU TOIT-NEUJMIN-DELPORTE Z. Sekanina, Jet Propulsion Laboratory, writes: “Very preliminary

analysis of the relative astrometry of the two brightest nuclei (cf. MPEC 2002-O10) has been completed, employing a new computer code recently developed by P. W. Chodas and myself. The parameters of the standard model for split comets are now determined with full account of the planetary perturbations and the nongravitational effects on the principal nucleus. The results suggest that nucleus B could have broken off from primary A near perihelion in 1996. If the event had occurred exactly at perihelion, plausible values for the nongravitational deceleration (4–8 × 10^{−5} solar attraction, as B is obviously a persistent companion; cf. Sekanina 1982, Comets, ed. L. L. Wilkening, pp. 251–287) require that B separated from A with a reasonably low velocity, whose transverse component ranged from 0.5 to 1 m/s in the direction opposite the orbital motion and whose normal component was some 0.4–0.5 m/s toward the north orbital pole. These solutions are independent of the radial component of the separation velocity. Similar solutions are also found for separation times 100 days before and after perihelion, except that the deceleration then correlates with both the transverse and radial components. Because of the comet’s extremely low orbital inclination, it is doubtful that the separation parameters can ever be determined with high accuracy. All examined solutions yield essentially the same ephemeris, which shows that the projected separation of B from A will diminish in the coming weeks. The offsets and position angles are as follows (0^{h} UT): 2002 Aug. 4, 853′′, 259^{∘}.1; 14, 814′′, 259^{∘}.1; 24, 758′′, 259^{∘}.0; Sept. 3, 694′′, 258^{∘}.7; 13, 631′′, 258^{∘}.3; 23, 571′′, 257^{∘}.7; Oct. 3, 516′′, 257^{∘}.0. It is unlikely that companions C–T (cf. IAUC 7935) are all products of the same event. In particular, C–F were probably released from A more recently than B was. Some of nuclei M–T may be fragments of B, but a more complex fragmentation hierarchy is also possible. Accurate astrometry on existing images and additional observations may allow one to make more, but not very, definite statements in the future.”

SUPERNOVA 2002ec IN NGC 5910 M. Salvo and P. Price, Research School of Astronomy and Astro-

physics, Australian National University (ANU), report that a spectrum (range 340–920 nm, resolution ∼ 0.4 nm) of SN 2002ec (cf. IAUC 7937), taken with the ANU 2.3-m telescope (+ double beam spectrograph) at Siding Spring Observatory on July 31.43 UT, shows that this object is a Type Ia supernova close to maximum light.

2002 August 3 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother